Brad drives 450km at 80km/hr. He takes 3-15 minutes breaks. How much time does his trip take in hours?
Answer by John10(297)

Brad drives 450km at 80km/hr. He takes 3-15 minutes breaks. How much time does his trip take in hours?
Time = distance/speed = 450/80 = 5.625 hours
he takes 45 minutes = 0.75 hour break. so the total trip is 5.625 - 0.75 = 4.875 hours
John10:)